OpenStudy (skyepurple1997):

This is talking about disjointed/ mutually exclusive. (They can't happened at the same time.) One card is randomly selected from a stranded deck. Compute the probability A. Of randomly selecting a two or three B. Of randomly selecting a two or club from a deck of cards. C.not selecting a heart or a face card.

OpenStudy (amistre64):

your thoughts?

OpenStudy (amistre64):

i got no idea what a stranded deck is ...

OpenStudy (amistre64):

standard maybe?

OpenStudy (skyepurple1997):

yes, sorry.

OpenStudy (amistre64):

2 or a 3 ... how many cards in the deck have a 2 or a 3 on them?

OpenStudy (amistre64):

the mathical formula would be: P(A) + P(B) - P(AnB)

OpenStudy (skyepurple1997):

there are 4 2's and 4 3's

OpenStudy (amistre64):

good, so 8 cards that we can pull that are either a 2 or a 3 8 out of 52 right? P(2) = 4/52 P(3) = 4/52 P(2 n 3) = 0/52 so the formula works fine as well

OpenStudy (amistre64):

if we remove all the 2s and all the clubs from the deck, how many cards do we have?

satellite73 (satellite73):

just to butt in, it is "\(\text{disjoint"}\) not \(\text{disjoint}\color{red}{ed}\)

OpenStudy (amistre64):

well, maybe a stranded deck of cards IS disjointed lol

OpenStudy (skyepurple1997):

what does P(2 n 3) = 0/52 suppose to mean? And yes lol i know, sorry for the typos lol

OpenStudy (amistre64):

the probability that a card is both a 2 and a 3 is, 0 out of 52 cards.

OpenStudy (skyepurple1997):

ook, i understand. it just looked weird lol

OpenStudy (amistre64):

math is weird, yes

OpenStudy (skyepurple1997):

p(2 or 3) = 4/52 + 4/52

OpenStudy (amistre64):

yes ... but for completness, we would want to excluded double counting so we remove the set that is common to both 2 and 3 ... otherwise we count them twice.

OpenStudy (skyepurple1997):

8/52 = 4/26 = 2/13 =0.1539 is that right?

OpenStudy (skyepurple1997):

oh, what? sooooo..

OpenStudy (amistre64):

consider the venn:  n(A or B) = n(A) + n(B) = (1+2)+(2+3) we have something counted twice, we have to remove it ... n(AnB) n(A or B) = n(A) + n(B) - n(AnB) = (1+2)+(2+3) - (2)

OpenStudy (skyepurple1997):

how would you know something is counted twice?

OpenStudy (amistre64):

the other questions make this more obvious ...

OpenStudy (amistre64):

how many cards have a 2 on them? how many cards are clubs? if we count all the 2s with all the clubs .... are we overcounting something?

OpenStudy (skyepurple1997):

yes

OpenStudy (skyepurple1997):

this is consider a dependent situation, right?

OpenStudy (amistre64):

dependence is a specific concept in probability ... it is easily confused with mutually exclusive as a word, but it is defined as something entirely different. if the probability of an event is the same for any given event that it occurs in .... then the probability of the event does not depend on a given situation. this does not mean that the events are or are not mutually exclusive ... its just that the probability works out so that it is the same regardless.

OpenStudy (amistre64):

spose that 1/4 of the m&ms in a bowl are green. spose that after eating 12 m&ms, we determine that 1/4 of the remaining m&ms are still green. in this instance, the probability remains the same regardless of which event we observe. so we say conclude that it simply doesnt matter which events probability we use.

OpenStudy (amistre64):

P(green)= P(green, given 12 m&ms are missing) 1/4 = 1/4 the probability of a green m&m is the same regardless of the event we choose ... the probability is independant of the event that you calculate it for.

OpenStudy (skyepurple1997):

ok, i think i get what you're saying.

OpenStudy (amistre64):

say 1/3 of the class has blond hair and given that we pull out all the girls, we find that 1/3 of the girls have blond hair. the probability of picking a blond is the same for each event, the probability does not depend on choosing from the class or from the subset.

OpenStudy (skyepurple1997):

ok.

OpenStudy (skyepurple1997):

Its starting to make sense to me

OpenStudy (amistre64):

just drilling in that words used in math have more precise definition then assuming they are just english words

OpenStudy (amistre64):

Mathical terminology should not have 'synonyms' English usage and math usage should be compartmentalized in your head :)

OpenStudy (skyepurple1997):

I actually agree

OpenStudy (amistre64):

so number of 2s in a deck: 2,2,2,2 number of clubs in a deck: A,2,3,4,5,6,7,8,9,J,Q,K 4+13 is not a good count is it? something has been counted more than once in this case. a single card is both a 2 and a club ... we counted it in both sets, we need to remove the duplication. n(2) = 4 n(clubs)=13 n(2 n clubs)=1 4+13-1 is a proper count for the desired outcomes right?

OpenStudy (skyepurple1997):

yes, i agree

OpenStudy (skyepurple1997):

so, it would be 4/52 + 13/52 = 17/52 - 1/52 = 16/52

OpenStudy (amistre64):

correct Of randomly selecting a two or club from a deck of cards. (4+13-1) out of (52) is our results

OpenStudy (amistre64):

now the last questions is: not selecting a heart or a face card how many hearts? how many face cards? how many that are both heart and face?

OpenStudy (skyepurple1997):

c. is of selecting, not, not selecting. that was a typo real question: of selecting a heart or a face card

OpenStudy (skyepurple1997):

so, in this situation, i would need to find how many hearts and how many faces, right?